OpenStudy (anonymous):

a-3a+c+4a-2b+6-a+b-5c

OpenStudy (anonymous):

Combine like terms.

OpenStudy (anonymous):

If \(ac+bc\), then \((a+b)c\).

OpenStudy (anonymous):

i know but idont knoq how to add those 2 a's

OpenStudy (anonymous):

add/subtract "LIKE" terms only -2a + c + 4a - 2b + 6 + a + b - 5c 2a + c - 2b + 6 + a + b - 5c 3a + c - 2b + 6 + b - 5c 3a -4c -b + 6

OpenStudy (anonymous):

You have \(a-3a+4a-a\). That is the same as \(1a-3a+4a-1a\). So: \((1-3+4-1)a\).

OpenStudy (unklerhaukus):

\[a-3a+c+4a-2b+6-a+b-5c\]\[=(a-3a)+(c)+(4a)+(-2b)+6+(-a)+b+(-5c)\]\[=(a-3a+4a-a)+(-2b+b)+(c-5c)+6\]\[=(1-3+4-1)a+(-2+1)b+(1-5)c+6\]